### v1

# Approximate Bayesian Computation sequential Monte Carlo via random forests

Abstract
Approximate Bayesian Computation (ABC) is a popular inference method when likelihoods are hard to come by. Practical bottlenecks of ABC applications include selecting statistics that summarize the data without losing too much information or introducing uncertainty, and choosing distance functions and tolerance thresholds that balance accuracy and computational efficiency. Recent studies have shown that ABC methods using random forest (RF) methodology perform well while circumventing many of ABC’s drawbacks. However, RF construction is computationally expensive for large numbers of trees and model simulations, and there can be high uncertainty in the posterior if the prior distribution is uninformative. Here we adapt distributional random forests to the ABC setting, and introduce Approximate Bayesian Computation sequential Monte Carlo with random forests (ABC-SMC-(D)RF). This updates the prior distribution iteratively to focus on the most likely regions in the parameter space. We show that ABC-SMC-(D)RF can accurately infer posterior distributions for a wide range of deterministic and stochastic models in different scientific areas.
1 Introduction
Mathematical modeling has played an important role in studying scientific phenomena. Its practical applications often depend on accurately extracting model parameters $\theta$ from experimentally observed data, $y^{obs}$. In the Bayesian framework, this entails inferring the posterior distribution
$\pi(\theta\mid y^{obs})\propto f(y^{obs}\mid\theta)\cdot\pi(\theta)$
of the parameters from the data, exploiting the likelihood $f(y^{obs}\mid\theta)$ of such observations under the model with given parameter values. However, the likelihood function is often intractable to derive theoretically, difficult to compute numerically, or is too complex to optimize directly. Approximate Bayesian Computation (ABC) was proposed as an alternative method to approximate the posterior distributions in such scenarios (Tavaré et al., 1997; Fu and Li, 1997; Pritchard et al., 1999; Beaumont et al., 2002). Using statistics to summarize model simulations, ABC seeks the parameters that result in minimal distance between the statistics and those of the observed data. Different variations have been proposed to improve the performance of the original ABC method, such as ABC Monte Carlo Markov Chain (Marjoram et al., 2003) and ABC Sequential Monte Carlo (Toni et al., 2008), and they have collectively found remarkable success in applications across different scientific areas (Sisson et al., 2018).
The accuracy and efficiency of ABC in practice depend on several factors. First, it requires a distance function to compare observed and simulated data. The metric is typically weighted so that different statistics contribute equally (Prangle, 2017). However, an equally important consideration is the relevance of each statistic, defined as the amount of information that it carries toward identifying the underlying parameters. Optimizing these two criteria can be challenging (Jung and Marjoram, 2011). Second, ABC requires a tolerance threshold to decide whether proposed parameters are accepted or rejected. The tolerance $\epsilon$ poses a trade-off between the computational efficiency, which increases with higher $\epsilon$, and accuracy, which improves as $\epsilon$ decreases, of the empirical posterior distributions. Choosing $\epsilon$ depends on the specific model, statistic choices and distance function, and it may require intuition or experimentation (Jung and Marjoram, 2011). Finally, and most importantly, ABC’s results depend on the choice of summary statistics. The approximated posterior distribution is only guaranteed to converge to the likelihood-based distribution for small $\epsilon$ if the statistics are sufficient. For complex models, it is not always possible or realistic to find low-dimensional sufficient statistics (Sisson et al., 2018, Chapter 5). This can sometimes be remedied by inclusion of many distinct statistics. However, the posterior estimation may become distorted if the selected statistics are noisy or uninformative. These factors in combination mean that optimizing the performance of ABC in practice can be challenging and require extensive experimentation.
Recent studies have shown that random forests (RF), a powerful non-parametric regression method (Breiman, 2001), can be employed in the ABC context to infer posterior distributions (Raynal et al., 2019). Its appeal lies in reduced dependence on user-defined hyperparameters that are essential to traditional ABC implementations, including the metric function and tolerance threshold. Moreover, RF has been shown to perform well even if the majority of statistics are pure noise, indicating that it is significantly more tolerant to a wider selection of statistics to represent the data (Raynal et al., 2019).
Therefore, RF methods have the potential to approximate the posterior distributions well with lower computational cost. However, the inclusion of all simulations in inferring parameters, coupled with wide range in the prior distributions for some models, can sometimes result in higher uncertainty in the posterior distributions.
In this paper, we propose a new inference method, Approximate Bayesian Computation sequential Monte Carlo with random forests (ABC-SMC-(D)RF). It inherits the non-parametric nature of RF methods, but is embedded in the framework of Sequential Monte Carlo (Toni et al., 2008). The posterior distribution evolves with successive iterations to focus on the most likely regions in the parameter space, resulting in more relevant model simulations and lower uncertainty in the final distributions.
We review some traditional ABC methods in Appendix A, and use them in comparisons with ABC-SMC-(D)RF described in the following sections. Recent ABC implementations using random forests are reviewed in Section 2, with illustrative examples. In Section 4, we describe ABC-SMC-(D)RF’s methodology. We then demonstrate its performance, compared both to traditional ABC algorithms and previous RF-based methods, in a wide range of deterministic and stochastic models (Section 5). We conclude with a discussion.
2 Random forests in the context of ABC
Although simple to implement, the ABC methods described in Appendix A depend heavily on the choices of hyperparameters, including the summary statistic function $S(y)$, the distance function $d(s,s^{\prime})$ and the tolerance threshold $\epsilon$ (Sisson et al., 2018). Furthermore, the choices of Markov kernels to propose new particles in ABC-MCMC and ABC-SMC also affect the approximated posterior distributions. As random forests (RF) gain in popularity as a powerful non-parametric regression technique (Breiman, 2001; Segal, 2004; Rigatti, 2017; Desai and Ouarda, 2021), its applications in the context of ABC are also becoming more prominent. RF-based methods eliminate the need for choice of metric, tolerance level and the perturbation kernels, making them an attractive alternative to the previous ABC implementations. Importantly, the RF method is relatively robust to noise (Marin et al., 2018), in that $S(y)$ can include many poorly informative statistics without significant impact on the results.
2.1 One-dimensional ABC random forests
Raynal et al. introduced ABC random forest (ABC-RF) to approximate the posterior distribution for ABC inference problems with one parameter (Raynal et al., 2019), based on the random forest formulation (Algorithm 1) developed by Breiman (2001). The root node for each tree $T_{t}$ is the reference table, consisting of parameters drawn from the prior distribution and simulated statistics. Each node $R_{*}$ in the decision tree is split into two nodes $R_{*1}$ and $R_{*2}$ by selecting one statistic $k_{*}\in\{1,\dots,|S|\}$, where $|S|$ is the total number of summary statistics, and a threshold $s_{*}$ such that the $L^{2}$-loss between the divided sets
$\frac{1}{|R_{*}|}\left(\sum_{(\theta,s)\in R_{*1}}\left(\theta-\bar{\theta}_{1}\right)^{2}+\sum_{(\theta,s)\in R_{*2}}\left(\theta-\bar{\theta}_{2}\right)^{2}\right)$ (1)
is minimized, where $|R_{*}|$ is the number of particles in $R_{*}$, and $\bar{\theta}_{1}$, $\bar{\theta}_{2}$ are the parameter means in $R_{*1}$ and $R_{*2}$ respectively. The tree is finished when each leaf either consists of less than $N_{min}$ particles, or all of the leaf’s particles have the same statistics. To guarantee forest diversity, the root node of each tree (Step 5 in Algorithm 1) is bootstrapped from the reference table, and the statistics considered for splitting each node (Step 9 in Algorithm 1) is limited to a randomly selected subset of size $n_{try}$ among the available statistics $S$.
1 Sample $\theta^{(1)},\dots,\theta^{(N)}\sim\pi(\theta)$
2Simulate $y^{(1)},\dots,y^{(N)}$ and compute $s^{(1)},\dots,s^{(N)}$
3Form reference table $R=\left\{\left(\theta^{(1)},s^{(1)}\right),\dots,\left(\theta^{(N)},s^{(N)}\right)\right\}$
4for $t=1,\dots,B$ do
5       $R_{0}\leftarrow$ bootstrapped from $R$
6      Tree $T_{t}$ is rooted in node $R_{0}$
7      while $\max_{\mathrm{leaf}R_{*}\in T_{t}}|R_{*}|>N_{min}$ do
8             Find a leaf $R_{*}$ in $T_{t}$ with $\left|R_{*}\right|>N_{min}$
9            Statistic candidates for splitting $R_{*}\leftarrow$ sample of size $n_{try}$ among $S$
10            Choose statistic $k_{*}$ among the candidates and threshold $s_{*}\in\mathbb{R}$ to split node $R_{*}$ into new leaves $R_{*1}=\left\{(\theta,s)\in R_{*}:s_{k_{*}}\leq s_{*}\right\}$ and $R_{*2}=\left\{(\theta,s)\in R_{*}:s_{k_{*}}>s_{*}\right\}$, such that the $L^{2}$-loss criterion (Eq. (1)) is minimized
11
Algorithm 1 Growing regression trees for ABC-RF (Raynal et al., 2019)
Given an observation $y^{obs}$, predicting its parameter $\theta$ based on tree $T_{t}$ involves following from the root node with the statistics $s^{obs}$ and comparing $s^{obs}_{k_{*}}$ with $s_{*}$ at each node $R_{*}$ to select the next node, until reaching a leaf $L_{t}(s^{obs})$. The prediction for $\theta$ based on $T_{t}$ is then the average parameter among particles in $L_{t}(s^{obs})$:
$\frac{\sum_{i=1}^{N}\theta^{(i)}\cdot n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}$
where $n_{t}^{(i)}$ is the number of times $\theta^{(i)}$ is duplicated in the bootstrapped sample $R_{0}$, and $\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}$ is the indicator for whether $s^{(i)}$ falls into the same leaf as $s^{obs}$. The prediction for $\theta$ based on the whole forest is the average of predictions based on each tree:
$\sum_{i=1}^{N}\theta^{(i)}\cdot\frac{1}{B}\sum_{t=1}^{B}\frac{n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}$
Raynal et al. (2019) argued that this weighted estimate implies that the density of particles $\theta^{(i)}$ with corresponding weights $w^{(i)}$ (Algorithm 2) forms the approximation $\pi_{ABC-RF}\left(\theta\middle|s^{obs}\right)$ for the posterior distribution.
1 for $t=1,\dots,B$ do
2       Follow the tree $T_{t}$ from root node with $s^{obs}$ until locating leaf $L_{t}(s^{obs})$
3for $i=1,\dots,N$ do
4       Weight for particle $\theta^{(i)}\leftarrow w^{(i)}=\frac{1}{B}\sum_{t=1}^{B}\frac{n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}{\sum_{i=1}^{N}n_{t}^{(i)}\cdot\textbf{1}_{s^{(i)}\in L_{t}(s^{obs})}}$
Algorithm 2 Posterior distribution from ABC-RF (Raynal et al., 2019)
Marin et al. (2022) developed an R package abcrf based on the ABC-RF method. For multivariate problems, ABC-RF is usually applied for each parameter marginally (Raynal et al., 2019). The authors also provided suggested values for RF hyperparameters; by default, $N_{min}=5$ and $n_{try}=|S|/3$. The most important hyperparameter is the count of trees $B$, for which they recommend analyzing whether the out-of-bag mean squared error stabilizes around the selected $B$ (Raynal et al., 2019; Pudlo et al., 2016).
2.2 An example from population genetics
We examine the performance of ABC-RF with a classic example from population genetics. We consider a sample of $n=1000$ DNA sequences collected from a stationary population that evolved according to the neutral coalescent model of Kingman (1982) with the infinitely-many-sites mutation model (ISM) (Kimura, 1969) with mutation rate $\theta$. Two statistics that are often used to summarize the mutational data are the number of segregating sites, $C_{n}$ and the site frequency spectrum (SFS) $f_{n}:=(f(1),\dots,f(n-1))$, where $f(j)$ is the number of mutations detected in exactly $j$ sequences. We have $C_{n}=\sum_{j=1}^{n-1}f(j)$, and
[…]
The distribution of $C_{n}$ is known explicitly (Tavaré, 1984, eqn. (9.5)), but for the present purposes it is enough to note that $f(j)$ has approximately a Poisson distribution with mean $\theta/j$ (Dahmer and Kersting, 2015), and that $C_{n}$ has approximately a Poisson distribution with mean given in (2). The likelihood of observing $C_{n}=c$ given $\theta$ and $n$ is therefore approximately the probability that $\text{Poisson}\left(\theta\sum_{i=1}^{n-1}\frac{1}{i}\right)=c$. If the prior distribution for $\theta$ is $\pi(\theta)\sim$ Uniform$\left(a,b\right)$, the approximate posterior distribution of $\theta$ is
$\pi(\theta\mid c)\propto\mathbb{P}\left[\text{Poisson}\left(\theta\sum_{i=1}^{n-1}\frac{1}{i}\right)=c\right],\quad a<\theta<b.$ (3)
We compare the posterior distributions approximated with ABC-RF (package abcrf) and ABC-REJ (package EasyABC, (Jabot et al., 2023)) with uniform prior distribution U$\left(1,20\right)$ and summary statistics $S=C$. The observed data consists of one simulation with $\theta^{obs}\sim$ U$\left(1,20\right)$, resulting in $C^{obs}=34$ segregating sites. We configure ABC-REJ and ABC-RF to employ $N=$ 10,000 simulations each, from which ABC-REJ retains the 500 simulations with values closest to $C^{obs}$, and ABC-RF infers $\theta$ from $B=500$ trees. The posterior distributions from ABC-REJ (median = 4.65, mean = 4.82, variance = 1.46) and ABC-RF (median = 4.82, mean = 4.91, variance = 1.47) are close to the true distribution (Eq. 3; median = 4.58, mean = 4.68, variance = 0.67) (Figure 1a).
We next analyze the results when $S=\left\{C_{n},f(1),f(2),\dots,f(m)\right\}$ summarize the sequenced mutations; we chose $m=\lfloor\sqrt{n}\rfloor=31$. The individual $f(j)$ values contain significantly more noise and less information, especially for larger $j$. Implementation of ABC-REJ is therefore challenging, as the distance function $d(s^{*},s^{obs})$ must be calibrated to put more weight on $C_{n}$ and less on the $f(j)$, in order to avoid an extremely low acceptance rate. The lack of metrics in ABC-RF means that it does not face this issue, and its approximation for the posterior distribution (median = 4.29, mean = 4.31, variance = 0.89) is largely identical to the case where $S=C_{n}$ (Figure 1b). Moreover, abcrf’s variable importance analysis confirms that $C_{n}$ is the most important statistic in the inference, followed by the $f(j)$ in essentially increasing order of $j$ (Figure 1d).
We perform a final inference experiment using $S=\left\{f(1),f(2),\dots,f(m)\right\}$ as the observed statistics. Although ABC-RF’s inferred posterior distribution for $\theta$ has higher variance compared to the previous two cases, its center is still close to the true value (Figure 1c, median = 4.30, mean = 4.59, variance = 2.37).
These examples highlight the potential of random forests in the context of ABC inference. ABC-RF does not require calibrating the distance functions and tolerance thresholds, which have significant impact on the efficiency and accuracy in inferring the posterior distribution with other ABC methods. Furthermore, the robustness to noise in simulation statistics indicates reduced significance of selecting which statistics to summarize the observed data.
Figure 1: Inference of mutation rate in the coalescent model. a, b, c: Posterior distributions inferred with ABC-REJ and ABC-RF for mutation rate $\theta$, compared against the true posterior distribution (Eq. (3)). The statistics used in the inference consist of the mutation count $C$ (a), $C$ and the SFS $f(j)$ with $j=1,\dots,31$ (b) or SFS $f(j)$ with $j=1,\dots,31$ (c). d: ABC-RF’s variable importance analysis for summary statistics in the inference problem in b.
2.3 Distributional random forests
Ćevid et al. (2022) developed distributional random forests (DRF) for multivariate regression problems. The approach considers two criteria to split tree nodes. The first criterion extends the $L^{2}$-loss formula (Eq. 1), which the authors rewrite as
$\frac{1}{|R_{*}|}\sum_{(\theta,s)\in R_{*}}\left(\theta-\bar{\theta}\right)^{2}-\frac{|R_{*1}|\cdot|R_{*2}|}{|R_{*}|^{2}}\left(\bar{\theta}_{2}-\bar{\theta}_{1}\right)^{2}$
where $\bar{\theta}$ is the parameter mean in the parent node $R_{*}$. Because the first term does not depend on the split, minimizing the $L^{2}$-loss is equivalent to maximizing
$\frac{|R_{*1}|\cdot|R_{*2}|}{|R_{*}|^{2}}\left(\bar{\theta}_{2}-\bar{\theta}_{1}\right)^{2}$ (4)
Choosing $(k_{*},s_{*})$ for multivariate problems then involves maximizing Eq. 4, aggregated across different parameters. Because this condition, denoted as CART criterion, only considers the difference in mean parameters in the child nodes and not the whole distributions, Ćevid et al. developed a second criterion based on the maximal mean discrepancy (MMD) metric (Ćevid et al., 2022). The MMD defines the difference between distributions of particles in two different sets by employing a positive-definite kernel and its embedding into a Reproducing Kernel Hilbert Space (RKHS) (Gretton et al., 2007). Due to the computational expense of MMD and the need to compute it for many different statistic and threshold candidates, the authors replace the kernel with its Fourier approximation. By default, they use the Gaussian kernel $k(\theta,\theta^{\prime})=\frac{1}{\left(\sigma\sqrt{2\pi}\right)^{|\Theta|}}\exp\left(\frac{-\lVert\theta-\theta^{\prime}\rVert_{2}^{2}}{2\sigma^{2}}\right)$ with parameter count $|\Theta|$ and bandwidth $\sigma$, the median pairwise $L^{2}$ distance between the parameters of the particles in the parent node. With Fourier features $\omega_{1},\dots,\omega_{L}$ randomly chosen from the multivariate Gaussian distribution $\text{Normal}_{|\Theta|}\left(\textbf{0},\sigma^{-2}\textbf{I}_{|\Theta|}\right)$ where 0 and $\textbf{I}_{|\Theta|}$ are square zero and identity matrices of size $|\Theta|\times|\Theta|$, Ćevid et al. derive the approximate MMD as
$\mathcal{D}(R_{*1},R_{*2})=\frac{1}{L}\sum_{l=1}^{L}\frac{|R_{*1}|\cdot|R_{*2}|}{|R_{*}|^{2}}\left|\frac{1}{|R_{*1}|}\sum_{\theta\in R_{*1}}\varphi_{\omega_{l}}(\theta)-\frac{1}{|R_{*2}|}\sum_{\theta\in R_{*2}}\varphi_{\omega_{l}}(\theta)\right|^{2}$ (5)
where $\varphi_{\omega}(u)=e^{i\omega^{T}u}$ are the approximate kernels. Constructing the random forest using the MMD criterion (Algorithm 3) consists of splitting each node, such that (5) is maximized.
1 Form reference table $R=\left\{\left(\theta^{(1)},s^{(1)}\right),\dots,\left(\theta^{(N)},s^{(N)}\right)\right\}$, similar to Algorithm 1
2for $t=1,\dots,B$ do
3       $R_{t}\leftarrow$ subsample of $R$ of size $n_{sample}$; $R_{t}$ is split into $R_{t1}$ and $R_{t2}$
4      Tree $T_{t}$ is rooted in node $R_{t1}$
5      while $\max_{\mathrm{leaf}R_{*}\in T_{t}}|R_{*}|>N_{min}$ do
6             Find a leaf $R_{*}$ in $T_{t}$ with $\left|R_{*}\right|>N_{min}$
7            Statistic candidates $\leftarrow$ subsample of $S$ with size $\min(\max(\tilde{n}_{try},1),|S|)$, where $\tilde{n}_{try}\sim\text{Poisson}(n_{try})$
8            if $\mathrm{criterion}=\mathrm{CART}$ then
9                   Choose statistic $k_{*}$ and threshold $s_{*}$ to split node $R_{*}$ into new leaves $R_{*1}=\left\{(\theta,s)\in R_{*}:s_{k_{*}}\leq s_{*}\right\}$ and $R_{*2}=\left\{(\theta,s)\in R_{*}:s_{k_{*}}>s_{*}\right\}$, such that the CART criterion (Eq. 4), aggregated across all parameters, is maximized
10            else if $\mathrm{criterion}=\mathrm{MMD}$ then
11                   $\sigma\leftarrow$ median pairwise distance between $\theta$’s in $R_{*}$
12                  Fourier features $\omega_{1},\dots,\omega_{L}\sim\text{Normal}_{|S|}\left(\textbf{0},\sigma^{-2}\textbf{I}_{|S|}\right)$
13                  Choose $k_{*}$ and $s_{*}$ to split $R_{*}$ into $R_{*1}$ and $R_{*2}$, such that the MMD criterion (Eq. (5)), defined with $\omega_{l}$’s, is maximized
14
15
16      for $i=1,\dots,\left|R_{t2}\right|$ do
17             Follow $T_{t}$ from root node with $s^{(i)}$ to locate leaf $L_{t}\left(s^{(i)}\right)$
18
Algorithm 3 Growing regression trees for DRF (Ćevid et al., 2022)
Once the random forest is constructed, computing the posterior distribution for an observation $y^{obs}$ consists of finding the leaf $L_{t}\left(s^{obs}\right)$ in each tree $T_{t}$ that the observed statistics $s^{obs}$ falls into. Because the samples for tree construction are not bootstrappped, the weight for each particle $\theta^{(i)}$ in the reference table is simply the normalized number of times that it ends up in the same leaf as $s^{obs}$ across the entire forest (Algorithm 4).
The CDF of the joint posterior distribution $\pi_{ABC-DRF}\left(\theta\middle|s^{obs}\right)$ is then approximated as
$\mathbb{P}\left(\theta_{1}\leq x_{1},\dots,\theta_{|\Theta|}\leq x_{|\Theta|}\right)=\sum_{i=1}^{N}w^{(i)}\textbf{1}_{\theta^{(i)}_{1}\leq x_{1},\dots,\theta^{(i)}_{|\Theta|}\leq x_{|\Theta|}}$
1 for $t=1,\dots,B$ do
2       Follow the tree $T_{t}$ from root node with $s^{obs}$ until locating leaf $L_{t}(s^{obs})$
3for $i=1,\dots,N$ do
4       Weight for particle $\theta^{(i)}\leftarrow w^{(i)}=\frac{1}{B}\sum_{t=1}^{B}\frac{\textbf{1}_{L_{t}\left(s^{i}\right)=L_{t}\left(s^{obs}\right)}}{\left|L_{t}\left(s^{obs}\right)\right|}$
Algorithm 4 Multivariate distribution from DRF (Ćevid et al., 2022)
Ćevid et al. (2022) developed an R package drf based on Algorithms 3 and 4. Compared to abcrf, the authors employ different methods to maximize forest diversity (Algorithm 3). Instead of a fixed number of $n_{try}$ randomly selected statistics from which splitting condition is chosen, the number of candidate statistics in DRF varies between nodes, following $\min(\max(\tilde{n}_{try},1),|S|)$ where $\tilde{n}_{try}$ is sampled from $\text{Poisson}(n_{try})$. Furthermore, instead of bootstrapping the reference table, the samples $R_{t1}$ to form each tree’s root node are subsampled from $R$. After the tree is constructed, the particles actually stored in the leaves for computing posterior weights are from a disjoint subsample $R_{t2}$ of $R$, in order to minimize overfitting. By default, the total number of particles used for tree construction $R_{t1}$ and weights $R_{t2}$ is $N/2$.
In the next section we describe how DRF may be adapted for use in approximate Bayesian computation.
3 ABC-DRF
The main limitation of abcrf is that it is typically used for inference of one parameter at a time, resulting in difficulty studying the joint posterior of the elements of $\theta$. We noted in Dinh et al. (2024) that DRF may be exploited for use in ABC by constructing the reference table just as for ABC-RF, and then using drf to generate observations from the joint posterior. We give an illustration of this approach, called ABC-DRF, in the next section.
3.1 A hierachical Normal mean example
We illustrate ABC-DRF with an adaptation of an example in Raynal et al. (2019) discussed in Dinh et al. (2024). The model has
$y_{1},y_{2},\ldots,y_{n}\mid\theta_{1},\theta_{2}$ $\sim$ ${\rm Normal}(\theta_{1},\theta_{2})$ (6)
$\theta_{1}\mid\theta_{2}$ $\sim$ $\text{Normal}(0,\theta_{2})$
$\theta_{2}$ $\sim$ $\text{IG}(\alpha,\beta)$
where $\text{IG}(\alpha,\beta)$ denotes the inverse gamma distribution with shape $\alpha$ and rate $\beta$ and density
$f(x;\alpha,\beta)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}(1/x)^{\alpha+1}\exp(-\beta/x),\quad x>0.$
The observed data $y^{obs}$ consists of $n$ data points $y^{obs}_{1},\dots,y^{obs}_{n}$.
The joint posterior distribution of $(\theta_{1},\theta_{2})$ is determined by
$\theta_{2}\mid y^{obs}$ $\sim$ $\text{IG}\left(\frac{n}{2}+\alpha,\,B\right)$ (7)
$\theta_{1}\mid\theta_{2},y^{obs}$ $\sim$ $\text{Normal}\left(\frac{n\cdot\bar{y}^{obs}}{n+1},\frac{2\cdot\theta_{2}}{n+1}\right),$ (8)
where
$B=\frac{1}{2}\left((S^{obs})^{2}+2\beta+\frac{n\cdot(\bar{y}^{obs})^{2}}{n+1}\right),\ \bar{y}^{obs}=n^{-1}\sum_{i=1}^{n}y_{i}^{obs},\ S^{obs}=\sqrt{\sum_{i=1}^{n}\left(y_{i}^{obs}-\bar{y}^{obs}\right)^{2}}.$
The marginal posterior of $\theta_{1}$ is
$\theta_{1}\mid y^{obs}\sim\frac{n\cdot\bar{y}^{obs}}{n+1}+\sqrt{\frac{2B}{(n+1)(n+2\alpha)}}\cdot\text{T}_{n+2\alpha}$
where $\text{T}_{m}$ denotes the $t$-distribution with $m$ degrees of freedom. $\theta_{1}$ and $\theta_{2}$ are uncorrelated under the posterior.
To illustrate the behavior of ABC-DRF, we generated a test set $y^{obs}=(y_{1}^{obs},\ldots,y_{10}^{obs})$ from the model, and let $s^{obs}=S(y^{obs})$ consist of the 61 summary statistics described in Raynal et al. (2019). The first three statistics $s^{obs}_{1},s^{obs}_{2},s^{obs}_{3}$ are the mean, variance and median absolute deviation of $y^{obs}$. The next eight statistics $s^{obs}_{4},\dots,s^{obs}_{11}$ are sums and products of either two or all values of $s^{obs}_{1},s^{obs}_{2},s^{obs}_{3}$. The final 50 statistics are noise: $s^{obs}_{12},\dots,s^{obs}_{61}\sim$ Uniform$(0,1)$.
We first infer the posterior distribution from drf, following the example in Dinh et al. (2024). The reference table consists of $N=10,000$ entries, each of which results from sampling $\theta$ and $y_{1},\dots,y_{n}$ from (6), then computing $s$ similarly to $s^{obs}$. The algorithm drf then infers the joint posterior distribution for $\theta$, which we compare against the true posterior distribution.
Figure 2a shows the comparison for $\alpha=4,\beta=5$. The posterior distributions from ABC-DRF are in agreement with the true density. Furthermore, the variable importance analysis detects information in $s_{1},\dots,s_{11}$, which indeed contain signals for the distribution of $y^{obs}$ (Figure 2b). In contrast, the pure noise statistics $s_{12},\dots,s_{61}$ are deemed unimportant, as expected.
Compared to abcrf, drf’s ability to infer the joint distribution is crucial for inference problems where the parameters are known to be dependent. We will experiment with such models in Section 5.
Figure 2: Inference of $\theta=(\theta_{1},\theta_{2})$ in the hierarchical model, with $\alpha=4,\beta=5$. a: Joint posterior distributions for $\theta_{1}$ and $\theta_{2}$, inferred from ABC-DRF with CART splitting rule from $N=10,000$ simulations (density heatmap) and ground truth (red contours, sampled from Eqs. 7 and 8), with marginal distributions for each parameter from ABC-DRF (blue histogram) and ground truth (red histogram). b: Variable importance analysis for summary statistics from ABC-DRF, performed with $N=1,000$ simulations.
4 ABC Sequential Monte Carlo with random forests
Despite their promise as relatively simple ABC approaches, there are some drawbacks in using abcrf and drf in practice. First, their performance typically improves as the forest size increases. The number of splitting criterion computations in growing a forest of $B$ balanced trees from a reference table of size $N$, where each split considers $n_{try}$ statistics, is $\mathcal{O}(B\cdot n_{try}\cdot N\cdot\log N)$. For complex problems involving many statistics and requiring large reference tables, the forest can become too computationally expensive to construct. Second, similar to other ABC implementations, the accuracy of the estimated posterior distributions depends on the number of simulations $N$. However, because the number of splits $\approx N\cdot\log N$ in balanced trees and more in unbalanced trees, the application of abcrf or drf in difficult inference problems where many simulations are necessary can demand more memory than available resources.
To alleviate these problems, we propose ABC sequential Monte Carlo with random forests (ABC-SMC-(D)RF). The algorithm starts with a reference table consisting of $N_{1}$ simulations computed with $\theta\sim\pi(\theta)$, from which either ABC-RF or ABC-DRF is applied to derive particle weights conditioned on $s^{obs}$. A new group of $N_{2}$ parameters is sampled from the weighted particle set, then perturbed with a Markov kernel $K_{2}(\theta|\theta^{\prime})$, to simulate new statistics. This forms the reference table for another iteration of the random forest algorithms, and the algorithm continues until iteration $T$. The parameters sampled from the final random forest iteration constitute the approximate posterior distribution we seek.
We introduce two versions of ABC-SMC using random forests: ABC-SMC-RF and ABC-SMC-DRF. The first version (Algorithm 5) is built around ABC-RF, and is applicable for problems with one parameter, or finding marginal distributions for multivariate problems. For multivariate models, ABC-RF is performed in each iteration $t$ for each parameter $j=1,\dots,|\Theta|$, using a reference table $R_{j}\in\mathbb{R}^{N_{t}\times(|S|+1)}$ consisting of the simulated statistics and corresponding values for $\theta_{j}$. The resulting weights are then used to sample $\theta_{j}$’s in the next iteration.
1 for $t=1,\ldots,T$ do
2       for $i=1,\dots,N_{t}$ do
3             if $t=1$ then
4                  Sample $\theta_{t}^{(i)}\sim\pi(\theta)$
5            else
6                   for $j=1,\dots,|\Theta|$ do
7                         Sample $\theta^{*}_{j}$ from $\left\{\theta^{(i)}_{j,t-1}\right\}$ with weights $\left\{w^{(i)}_{j,t-1}\right\}$
8                  $\theta^{*}\leftarrow\left(\theta^{*}_{1},\dots,\theta^{*}_{|\Theta|}\right)$
9                  $\theta_{t}^{(i)}\sim K_{t}(\theta|\theta^{*})$
10                  If $\pi\left(\theta_{t}^{(i)}\right)=0$, then return to Step 6
11
12            Simulate data $y_{t}^{(i)}$ with $\theta_{t}^{(i)}$, and compute $s_{t}^{(i)}=S\left(y_{t}^{(i)}\right)$
13
14      for $j=1,\dots,|\Theta|$ do
15             Form $R_{j}=\left\{\left(\theta_{j,t}^{(1)},s_{t}^{(1)}\right),\dots,\left(\theta_{j,t}^{(N_{t})},s_{t}^{(N_{t})}\right)\right\}$
16            Perform ABC-RF (Algorithms 1 and 2) with reference table $R_{j}$ to compute weights $w_{j,t}^{(1)},\dots,w_{j,t}^{(N_{t})}$ for observed statistics $s^{obs}$
17
Algorithm 5 ABC-SMC-RF for single parameters
The second version, ABC-SMC-DRF (Algorithm 6), is designed for inference of multivariate parameters. In each iteration, ABC-DRF is applied to a reference table $R\in\mathbb{R}^{N_{t}\times(|S|+|\Theta|)}$ that combines the entire parameter sets and statistics from all simulations. The weights that ABC-DRF predicts for $s^{obs}$ influence the sampling for parameter vectors in the next iteration.
1 for $t=1,\ldots,T$ do
2       for $i=1,\dots,N_{t}$ do
3             if $t=1$ then
4                  Sample $\theta_{t}^{(i)}\sim\pi(\theta)$
5            else
6                   Sample $\theta^{*}$ from $\left\{\theta^{(i)}_{t-1}\right\}$ with weights $\left\{w^{(i)}_{t-1}\right\}$
7                  $\theta_{t}^{(i)}\sim K_{t}(\theta|\theta^{*})$
8                  If $\pi\left(\theta_{t}^{(i)}\right)=0$, then return to Step 6
9
10            Simulate data $y_{t}^{(i)}$ with $\theta_{t}^{(i)}$, and compute $s_{t}^{(i)}=S\left(y_{t}^{(i)}\right)$
11
12      Form $R=\left\{\left(\theta_{t}^{(1)},s_{t}^{(1)}\right),\dots,\left(\theta_{t}^{(N_{t})},s_{t}^{(N_{t})}\right)\right\}$
13      Perform DRF (Algorithms 3 and 4) with reference table $R$ to compute weights $w_{t}^{(1)},\dots,w_{t}^{(N_{t})}$ for observed statistics $s^{obs}$
Algorithm 6 ABC-SMC-DRF for multiple parameters
We have developed an R package abc-smc-rf based on Algorithms 5 and 6. It offers several practical advantages when compared to direct applications of abcrf and drf. First, constructing iterative random forests of size $B$ from reference tables of sizes $N_{1},\dots,N_{T}$ requires $\mathcal{O}\left(B\cdot\sum_{t=1}^{T}(N_{t}\cdot\log N_{t})\right)$ splits, assuming balanced trees. This requires less computational resources than growing a random forest based on a reference table of size $N=\sum_{i=1}^{T}N_{i}$, which contains $\mathcal{O}\left(B\cdot N\log N)\right)$ splits.
Therefore, with the same total number of model simulations, ABC-SMC-(D)RF requires less memory and computational runtime. Second, similar to the behavior of ABC-SMC compared to ABC-REJ, ABC-SMC-(D)RF is likely to converge to the posterior distribution faster than ABC-RF or ABC-DRF, because the parameter distributions are constantly updated to focus on regions in the parameter space $\Theta$ that best explain $s^{obs}$. We will test this hypothesis with some examples in the next section. Finally, unlike ABC-RF and ABC-DRF which require the full reference table before forest construction and parameter prediction, the approximated posterior distribution from each iteration of ABC-SMC-(D)RF can be compared against the previous iteration to assess whether it has converged, potentially further lowering computational expense.
5 Results
In this section, we compare the performance of ABC-SMC-(D)RF (Algorithm 5 or 6) across different inference problems, against ABC-REJ (Algorithm 7), MCMC (Algorithm 8 or 9), ABC-SMC (Algorithm 10), ABC-RF (Algorithms 1, 2) and ABC-DRF (Algorithms 3, 4). Unless specified otherwise, for ABC-RF, ABC-DRF and ABC-SMC-(D)RF, we use the default parameters in abcrf and drf for tree count $B$, candidate statistic count $n_{try}$, leaf size threshold $N_{min}$, root node subsample size $n_{sample}$, etc.
5.1 Deterministic Lotka-Volterra model
The Lotka-Volterra model describes the interaction dynamics between predators and prey (Lotka, 1925; Volterra, 1928). The deterministic model for the number of prey $x$ and predators $y$ is in the form of paired nonlinear differential equations:
$\frac{\mathrm{d}x}{\mathrm{d}t}$ $=ax-cxy$
$\frac{\mathrm{d}y}{\mathrm{d}t}$ $=bxy-dy$
where $a$ is the prey’s birth rate, $d$ is the predator’s death rate, and $c$ and $b$ are the prey’s death rate and predator’s birth rate due to predation, respectively. We fix $c=d=1$ and seek to infer $\theta=(a,b)$ from prior distribution $a,b\sim\text{Uniform}(-10,10)$. Similar to Toni et al. (2008), we solve the ODE system with initial condition $(x(0),y(0))=(1,0.5)$ with $\theta=(1,1)$, and sample $(x,y)$ at eight time points between $t=0$ and $t=15$, then add Normal$(0,0.5^{2})$ noise to each data point to form the observed statistics $s^{obs}=\left\{x_{1},y_{1},\dots,x_{8},y_{8}\right\}$.
We first implement ABC-SMC-(D)RF for multiple parameters. Algorithm 6 is applied with $T=4$ iterations, each with $N_{t}=5,000$ simulations in the reference table, and perturbation kernels $K_{t}\left(\theta\middle|\theta^{*}\right)=\theta^{*}+\text{Uniform}(-0.1,0.1)$. Figures 3a, b show the distributions of $a$ and $b$ after each iteration $t$ in abc-smc-rf. The distributions become more concentrated as $t$ progresses, but stay centered around the true values of $(a,b)$ (Table 1).
To evaluate abc-smc-rf’s performance, we compare the final posterior distributions with ABC-DRF and ABC-SMC. Algorithm drf is performed on a reference table of size $N=20,000$, to match the total number of simulations in abc-smc-rf. We use the R package EasyABC (Jabot et al., 2023) to implement ABC-SMC with $T=5$ iterations. EasyABC uses the distance $d\left(s,s^{obs}\right)=\sum_{i=1}^{8}\left[(x_{i}-x_{i}^{obs})^{2}+(y_{i}-y_{i}^{obs})^{2}\right]$ to compare simulated statistics $s$ against observation $s^{obs}$, and adaptive perturbation kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)=\text{Normal}\left(\theta^{\prime},2\cdot\text{var}\left(\theta_{t-1}\right)\right)$ proposed in Beaumont et al. (2009). Referencing Toni et al. (2008), we impose the series of tolerance thresholds $\epsilon_{1}=30.0,\epsilon_{2}=16.0,\epsilon_{3}=6.0,\epsilon_{4}=5.0,\epsilon_{5}=4.3$. To produce 1,000 accepted particles, EasyABC requires $N=56,850$ simulations.
Figures 3c, d present the posterior distributions approximated by the ABC implementations. The marginal distributions from ABC-SMC-DRF are as centered and concentrated around the true values for $(a,b)$ as ABC-SMC (Table 1). In contrast, even though the distributions from DRF are centered around the ground truth, the variances are consistently higher, indicating more uncertainty.
Figure 3: Parameter inference for the deterministic Lotka-Volterra model. a, b: Iterative posterior distributions for reaction rates $a$ (a) and $b$ (b) from ABC-SMC-RF. c, d: Marginal posterior distributions for $a$ (c) and $b$ (d), from ABC-SMC, ABC-DRF and ABC-SMC-DRF. Black vertical lines denote true parameter values.
Statistics ABC-SMC-DRF iterations ABC-DRF ABC-SMC
1 2 3 4 (Final posterior)
$\mathbb{E}(a)$ 0.8237 0.3349 1.2044 1.1215 0.7562 1.2912
Var(a) 1.9018 0.2757 0.1279 0.0333 0.5953 0.1104
$\mathbb{E}(b)$ 1.6816 1.4889 1.0061 0.9704 1.3092 1.0269
Var(b) 1.5298 0.1536 0.0697 0.0313 0.3593 0.1049
Table 1: Means and variances of marginal posterior distributions for the deterministic Lotka-Volterra model from ABC-SMC-DRF, ABC-DRF, and ABC-SMC. The best result for each statistic across different algorithms is in bold ($\mathbb{E}(a)$, $\mathbb{E}(b)$ closest to true values $(a,b)=(1,1)$, and lowest variance in each statistic).
5.2 Linear birth-death process
We next examine ABC-SMC-DRF’s performance for the linear birth-death process example in Tavaré (2018). The model starts at time 0 with $Z(0)$ individuals, each of which divides at rate $\lambda>0$ and dies at rate $\mu>0$. The number of individuals at time $t>0$ is denoted $Z(t)$, with expected value $\mathbb{E}Z(t)=Z(0)\mathrm{e}^{(\lambda-\mu)t}$.
The probability $\mathbb{P}\left(Z(t_{i+1})=z_{i+1}\middle|Z(t_{i})=z_{i},\lambda,\mu\right)$ that there are $z_{i+1}$ individuals at time $t_{i+1}$ given that there are $z_{i}$ individuals at time $t_{i}$ follows from Keiding (1975) as
$\sum_{l=0}^{\min(z_{i},z_{i+1})}\binom{z_{i}}{l}(1-\alpha(t))^{l}\alpha(t)^{z_{i}-l}\cdot\binom{z_{i+1}-1}{z_{i+1}-l}(1-\beta(t))^{l}\beta(t)^{z_{i+1}-l}$
where $\alpha(t)$ and $\beta(t)$ are defined by
[…]
(Kendall, 1948). The Markov property implies that the likelihood for $\lambda$ and $\mu$, given an observation $y^{obs}$ where $Z(t_{1})=z_{1},\dots,Z(t_{n})=z_{n}$, is:
$f\left(\lambda,\mu\middle|y^{obs}\right)=\prod_{i=0}^{n-1}\mathbb{P}\left(Z(t_{i+1})=z_{i+1}\middle|Z(t_{i})=z_{i},\lambda,\mu\right)$ (9)
Our inference problem consists of approximating the posterior distribution of $\theta=(\lambda,\mu)$ from a prior distribution $\lambda,\mu\sim$ Uniform$(0,20)$, for observed data $s^{obs}=y^{obs}=(z_{i},t_{i})_{i=1,\dots,25}$ from initial condition $Z(0)=10$ in (Tavaré, 2018, Table 3). The true posterior distribution can be computed from (9). Similarly, we build a MCMC chain using the Metropolis-Hastings algorithm 8 with likelihood function (9) and uniform proposal kernel $q\left(\lambda,\mu\middle|\lambda^{\prime},\mu^{\prime}\right)=(\lambda^{\prime}+\text{U}(-1,1),\mu^{\prime}+\text{U}(-1,1))$ to approximate the posterior distribution. After 2500 burn-in steps, we accepted 1950 simulations from a total of $N=100,000$ steps (average acceptance rate = 0.379) spaced every 50 steps. Finally, we implement drf with a reference table of size $N=20,000$, abcrf separately for $\lambda$ and $\mu$ likewise with $N=20,000$, and abc-smc-rf using the multivariate option with $T=4$ iterations, each with $N_{t}=5,000$ simulations. All RF methods approximate the posterior distributions from $B=2,500$ trees. ABC-SMC-DRF implements perturbation kernels $K_{t}(\theta|\theta^{\prime})=\theta^{\prime}+\text{U}(-2,2)\times\text{U}(-2,2)$.
Unsurprisingly, compared to the true density, MCMC is the most accurate in its inferred joint and marginal posterior distributions (Figure 4), as it utilizes the likelihood function in accepting new particles in the chain. The joint distributions from ABC-DRF and ABC-SMC-DRF are concentrated in the same area as MCMC, and exhibit a similar correlation pattern between $\lambda$ and $\mu$ (Figure 4a). However, quantile-quantile plots comparing the marginal distributions of $\lambda$ and $\mu$ between the exact posterior and the estimates reveal that ABC-DRF’s result has a significant bias toward larger values, compared to MCMC (Figure 4b, c). The marginal distributions from ABC-RF have reduced tails and are closer to the true density. However, the results from ABC-SMC-DRF are the most accurate among the RF implementations, and are almost as good as those from MCMC (Figure 4b, c).
Figure 4: Parameter inference for the linear birth-death branching process. a: Joint posterior distributions for birth rate $\lambda$ and death rate $\mu$, computed with the likelihood (Eq. 9, density heatmap) and estimated with MCMC, ABC-DRF with CART splitting rule, and ABC-SMC-DRF. b, c: Marginal quantile-quantile plots for inferred $\lambda$ (b) and $\mu$ (c) from MCMC, ABC-RF, ABC-DRF, and ABC-SMC-DRF (y axis), against the exact likelihood-based distributions (x axis).
5.3 Stochastic biochemical reaction systems
Finally, we examine ABC-SMC-DRF’s application in inferring reaction rates for biochemical systems. We first examine the Michaelis-Menten model, which describes the kinetics between an enzyme $E$ that binds to a substrate $S$ to form a complex $ES$, leading to product $P$:
$ES$ $\rightarrow E+P$ (10)
$E+S$ $\rightarrow ES$ (11)
$ES$ $\rightarrow E+S$ (12)
The state of the system at time $t$ can be described as $X(t)=\left[E(t),S(t),ES(t),P(t)\right]$, the counts of each type of molecule. The time for the next reaction (10), (11) and (12) to occur is exponentially distributed, with rates defined by their propensity functions:
$\alpha_{1}(X(t))$ $=\bar{c}_{1}\cdot ES(t);\quad\bar{c}_{1}=c_{1}$
$\alpha_{2}(X(t))$ $=\bar{c}_{2}\cdot E(t)\cdot S(t);\quad\bar{c}_{2}=\frac{10^{c_{2}}}{\text{nA}\cdot\text{vol}}$
$\alpha_{3}(X(t))$ $=\bar{c}_{3}\cdot ES(t);\quad\bar{c}_{3}=10^{c_{3}}$
respectively, where the Avagadro’s constant nA $=6.023\cdot 10^{23}$ approximates the number of molecules in a mole, vol $=10^{-15}$ is the volume of the system, and $\theta=(c_{1},c_{2},c_{3})$ parameterizes the Michaelis-Menten model for different biochemical systems (Wilkinson, 2018, Example 7.3).
The initial state is defined by $E(0)=2\cdot 10^{-7}\cdot\text{nA}\cdot\text{vol}$, $S(0)=5\cdot 10^{-7}\cdot\text{nA}\cdot\text{vol}$, and $ES(0)=P(0)=0$ (Wilkinson, 2018). We seek to infer $\theta$ from $s^{obs}=\left\{X(t),t=1,\dots,10\right\}$, simulated with true parameters $c_{1}=0.1,c_{2}=6,c_{3}=-4$. We compare the results from drf with $N=$ 20,000 simulations, and abc-smc-rf with $T=$ 5 iterations, each with reference tables of size $N_{t}=$ 4,000, from uniform prior distributions $\pi\left(c_{1}\right)=$ U$(0,1)$, $\pi\left(c_{2}\right)=$ U$(5,7)$, $\pi\left(c_{3}\right)=$ U$(-5,-3)$. The perturbation kernels in abc-smc-rf in each iteration are $K_{t}\left(c_{1}\middle|c_{1}^{\prime}\right)=c_{1}^{\prime}+\text{U}(-0.05,0.05)$, $K_{t}\left(c_{2}\middle|c_{2}^{\prime}\right)=c_{2}^{\prime}+\text{U}(-0.1,0.1)$, and $K_{t}\left(c_{3}\middle|c_{3}^{\prime}\right)=c_{3}^{\prime}+\text{U}(-0.1,0.1)$. All simulations are generated by applying Gillespie’s algorithm (Gillespie, 1977).
Comparing the results for $c_{1}$ and $c_{2}$ shows that ABC-SMC-DRF’s posterior distributions ($c_{1}$: mean = 0.110, 95% CI = $[0.083,0.141]$; $c_{2}$: mean = 5.959, 95% CI = $[5.905,6.016]$) are centered closely around the true values of $c_{1}=0.1,c_{2}=6$, similarly to ABC-DRF, but with narrower confidence intervals (ABC-DRF $c_{1}$: mean = 0.102, 95% CI = $[0.023,0.176]$; $c_{2}$: mean = 5.943, 95% CI = $[5.759,6.093]$) (Figure 5a, b). In contrast, both methods are uncertain about the true distribution for $c_{3}$ (ABC-SMC-DRF: mean = -3.994, 95% CI = $[-4.887,-3.259]$; ABC-DRF: mean = -3.950, 95% CI = $[-4.915,-3.032]$) (Figure 5c).
Analyzing the reaction system reveals a likely reason: assuming true values of $\theta$, reactions (10) and (12) occur with rates $\bar{c}_{1}=0.1$ and $\bar{c}_{3}=10^{-4}$, respectively. Because of the big difference in magnitudes, complex $ES$ overwhelmingly undergoes reaction (10), while reaction (12) rarely occurs and does not significantly impact the observed molecule counts.
We apply Morris’s global sensitivity analysis (Morris, 1991) to study the parameter identifiability (R package sensitivity (Iooss et al., 2024; Monari and Strachan, 2017)). The method involves computing the elementary effect of each parameter, defined in this case as the change in the statistics divided by the change in the parameter, across different sampling schemes, and analyzing its absolute mean $\mu^{*}$ and standard deviation $\sigma$. The study shows that $c_{1}$ ($\mu^{*}=0.019,\sigma=0.012$) and $c_{2}$ ($\mu^{*}=0.018,\sigma=0.002$) have significantly higher absolute mean elementary effects than $c_{3}$ ($\mu^{*}=0.002,\sigma=0.003$) (Figure 5d). This implies that varying $c_{3}$ has little to no effect on the observed statistics, consistent with earlier findings (Degasperi and Gilmore, 2008).
Indeed, despite the uncertainty in $c_{3}$, simulations performed with $\theta$-values sampled from the posterior distributions of ABC-DRF and ABC-SMC-DRF are centered around the observed values $s^{obs}$ across different molecules and time points (Figure 5e, f, g, h). Moreover, consistent with ABC-SMC-DRF’s higher certainty in the inference for $c_{1}$ and $c_{2}$, molecule counts simulated from its posterior distribution have significantly reduced range compared to ABC-DRF.
Figure 5: Parameter inference for the Michaelis-Menten reaction system. a, b, c: Marginal distributions for reaction rates $c_{1}$ (a), $c_{2}$ (b) and $c_{3}$ (c), inferred with ABC-DRF and ABC-SMC-DRF, compared against true values (black lines). d: Absolute mean ($\mu^{*}$) and standard deviation ($\sigma$) of elementary effects of $c_{1},c_{2}$ and $c_{3}$, using Morris’s global sensitivity analysis. e, f, g, h: Range of molecule numbers for $E$ (e), $S$ (f), $ES$ (g) and $P$ (h) across time, simulated with $\theta\sim\pi_{ABC-DRF}\left(\theta\middle|s^{obs}\right)$ and $\pi_{ABC-SMC-DRF}\left(\theta\middle|s^{obs}\right)$ (box plots), compared against observed data $s^{obs}$ (black dots and lines).
6 Discussion
In this paper, we introduce a new Bayesian inference method, Approximate Bayesian Computation sequential Monte Carlo with random forests (ABC-SMC-(D)RF). It inherits random forest’s non-parametric nature, resulting in less dependence on user-defined hyperparameters compared to traditional ABC methods such as ABC-REJ, MCMC and ABC-SMC. The random forest is embedded in a sequential Monte Carlo regime which progressively updates the parameter distributions to focus on regions in the parameter space with higher likelihood. We perform numerical experiments for deterministic and stochastic models in ecology, population genetics and systems biology, and observe that ABC-SMC-(D)RF results in better posterior approximations than previous RF methods, and comparable results with traditional ABC implementations with optimized hyperparameters.
However, there are several areas of improvement that can extend ABC-SMC-(D)RF’s capabilities. First, a hyperparameter in ABC-SMC-(D)RF that potentially has an impact on the results is the choice of perturbation kernels $K_{t}\left(\theta\middle|\theta^{\prime}\right)$, which should balance between exploring the parameter space and targeting the regions already found to contain high likelihood from previous iterations. In ABC-MCMC and ABC-SMC, it is common to choose uniform or Gaussian kernels, but the optimal kernel form and parameterization for specific problems may be complicated (Drovandi and Pettitt, 2011; Lee, 2012). A typical approach is to employ kernels that are adjusted dynamically depending on the performance of the previous iteration (Beaumont et al., 2009; Del Moral et al., 2012; Liu et al., 2000; Filippi et al., 2013; Atchadé and Fort, 2010). A potentially promising improvement for ABC-SMC-(D)RF would be to adopt a similar adaptive kernel framework.
Second, it is sometimes not necessary to continue until the final iteration in ABC-SMC-(D)RF, if further computation is unlikely to yield significant improvements in the posterior distributions. There have been a number of papers determining the stopping criterion in ABC-SMC (for example, Prangle (2017)), and their incorporation in ABC-SMC-(D)RF promises to lower the computational cost while retaining accuracy.
Finally, in this paper we have focused on parameter inference. ABC-RF can perform model selection Raynal et al. (2019), and the same framework could potentially be implemented in ABC-DRF and ABC-SMC-(D)RF. ABC-SMC-(D)RF is designed as a wrapper around abcrf (Marin et al., 2022) and drf (Michel and Ćevid, 2021) and can therefore be updated whenever the original libraries are updated.
Code availability
The code for ABC-SMC-(D)RF and studies performed in this paper are available at https://github.com/dinhngockhanh/abc-smc-rf.
Acknowledgments
The authors acknowledge the support from the Herbert and Florence Irving Institute for Cancer Dynamics and Department of Statistics at Columbia University.
Additional information
The authors declare no competing interests.
Appendix A Review of some ABC approaches
Assume we have a model $y=\mathcal{M}(\theta)$ that describes data $y$ with parameters $\theta\in\Theta$, and a prior distribution $\pi(\theta)$ that describes our a priori knowledge about the parameters. The posterior distribution of $\theta$ given observed data $y^{obs}$ is
$\pi(\theta|y^{obs})\propto f(y^{obs}|\theta)\cdot\pi(\theta),$
where $f(y^{obs}|\theta)$ is the likelihood function.
Approximate Bayesian Computation (ABC) was proposed as a practical method to approximate posterior distributions for problems in which the likelihood function is either theoretically intractable or computationally expensive. The data $y^{obs}$ is summarized with some low-dimensional statistic $s^{obs}=S(y^{obs})$. ABC methods then find the posterior distribution $\pi\left(\theta\mid d(s,s^{obs})<\epsilon\right)$, where $\epsilon$ is a tolerance threshold and $d(s,s^{obs})$ is a metric to compare $s^{obs}$ and statistics $s$ from data simulated with parameter $\theta$. If the statistic $S$ is sufficient, then $\pi\left(\theta\middle|d(s,s^{obs})<\epsilon\right)$ approximates the likelihood-based $\pi\left(\theta\middle|y^{obs}\right)$ as $\epsilon$ approaches 0. Here we briefly review some existing methods that will be used in comparisons with the proposed random forest approaches.
A.1 Rejection method
The simplest implementation of the ABC approach is the rejection method (ABC-REJ) (Algorithm 7), originally developed in population genetics [Tavaré et al., 1997, Fu and Li, 1997, Pritchard et al., 1999, Beaumont et al., 2002]. The statistics $s^{*}$ from data $y^{*}$ simulated with parameters $\theta^{*}\sim\pi(\theta)$ are compared to the observed statistics $s^{obs}$ with a distance function $d$, and only $\theta^{*}$ with $d(s^{*},s^{obs})$ smaller than a tolerance threshold $\epsilon$ are retained. The accepted values $\left\{\theta^{(1)},\dots,\theta^{(N)}\right\}$ are taken to approximate the required posterior distribution.
1 for $i=1,\dots,N$ do
2       Sample $\theta^{*}\sim\pi(\theta)$
3      Simulate data $y^{*}$ from the model with parameter $\theta^{*}$
4      Compute summary statistics $s^{*}=S(y^{*})$
5
6Order $\theta^{*}$ by increasing $d\left(s^{*},s^{obs}\right)$
Select $\theta^{(1)},...$ from the top percentiles of $\left\{\theta^{*}\right\}$
Algorithm 7 ABC-REJ
A.2 Markov Chain Monte Carlo
For problems where the prior distribution and posterior distributions are significantly different, the ABC-REJ method might require sampling $\theta^{*}$ and simulating $y^{*}$ many times before finding a parameter set that satisfies $d(s^{*},s^{obs})<\epsilon$ Different techniques have been proposed to improve the acceptance rate and decrease the number of model simulations.
Markov chain Monte Carlo (MCMC) approximates the posterior distribution by constructing a Markov chain, where the conditional distribution of the next proposed parameters only depends on the current set [Metropolis et al., 1953, Hastings, 1970]. Under suitable regularity conditions, the Markov chain converges to the posterior distribution[Tierney, 1994]. The Metropolis-Hastings algorithm 8 builds a Markov chain based on the likelihood function, starting from an initial $\theta^{(0)}$ sampled from the prior distribution [Metropolis et al., 1953, Hastings, 1970]. At each step, it proposes the next parameter set from a Markov kernel function $q\left(\theta\middle|\theta^{\prime}\right)$. The resulting chain $\left\{\theta^{(1)},\dots,\theta^{(N)}\right\}$ approximates the true posterior distribution $\pi\left(\theta\middle|y^{obs}\right)$.
1 Sample $\theta^{(0)}\sim\pi(\theta)$
2for $i=1,\dots,N$ do
3       Propose candidate parameter $\theta^{*}\sim q\left(\theta\middle|\theta^{(i-1)}\right)$
4      Calculate the acceptance probability $p_{accept}=\min\left(1,\frac{f\left(\theta^{*}\middle|y^{obs}\right)\cdot\pi\left(\theta^{*}\right)\cdot q\left(\theta^{(i-1)}\middle|\theta^{*}\right)}{f\left(\theta^{(i-1)}\middle|y^{obs}\right)\cdot\pi\left(\theta^{(i-1)}\right)\cdot q\left(\theta^{*}\middle|\theta^{(i-1)}\right)}\right)$ and random threshold $p\sim$ Uniform$(0,1)$
5      If $p_{accept}>p$, then $\theta^{(i)}\leftarrow\theta^{*}$, otherwise $\theta^{(i)}\leftarrow\theta^{(i-1)}$
Algorithm 8 Metropolis-Hastings MCMC
Marjoram et al. [2003] developed an ABC-MCMC method (Algorithm 9), a likelihood-free alternative to the Metropolis-Hastings algorithm. In lieu of $f\left(\theta\middle|y^{obs}\right)$, a proposed $\theta^{*}$ is only accepted if $d(s^{*},s^{obs})<\epsilon$. Because $\theta^{*}$ is drawn close to a previously accepted parameter set, the condition is typically more likely to be satisfied compared to the ABC-REJ method [Sisson et al., 2018]. The Markov chain results in the ABC-MCMC posterior distribution $\pi_{ABC-MCMC}\left(\theta\mid d(s^{*},s^{obs})<\epsilon\right)$.
1 Sample $\theta^{(0)}\sim\pi(\theta)$
2for $i=1,\dots,N$ do
3       Propose candidate parameter $\theta^{*}\sim q\left(\theta\middle|\theta^{(i-1)}\right)$
4      Simulate data $y^{*}$ with $\theta^{*}$, and compute summary statistics $s^{*}=S(y^{*})$
5      If $d(s^{*},s^{obs})\geq\epsilon$, then $\theta^{(i)}\leftarrow\theta^{(i-1)}$ and return to Step 3 with $i\leftarrow i+1$
6      Calculate probability to accept $p_{accept}=\min\left(1,\frac{\pi\left(\theta^{*}\right)\cdot q\left(\theta^{(i-1)}\middle|\theta^{*}\right)}{\pi\left(\theta^{(i-1)}\right)\cdot q\left(\theta^{*}\middle|\theta^{(i-1)}\right)}\right)$ and random threshold $p\sim$ Uniform$(0,1)$
7      If $p_{accept}>p$, then $\theta^{(i)}\leftarrow\theta^{*}$, otherwise $\theta^{(i)}\leftarrow\theta^{(i-1)}$
Algorithm 9 ABC-MCMC
To improve reliability and accuracy of the results, it is common practice to discard the particles in the initial “burn-in” phase of the Markov chain in both MCMC and ABC-MCMC, as these particles are dependent on the starting point and need not be representative of the posterior distribution.
A.3 Sequential Monte Carlo
Because the accepted particles in ABC-MCMC are correlated, the Markov chain can be stuck in regions of the parameter space with low probability. If the tolerance threshold $\epsilon$ is too small, the chain might take a long time to collect enough accepted particles [Sisson et al., 2007]. Moreover, the Markov chain cannot easily be computed in parallel.
Toni et al. [2008] introduced the ABC sequential Monte Carlo (ABC-SMC) method (Algorithm 10) as a different approach to solve the problem of low acceptance rate in ABC-REJ. ABC-SMC refines the particle population drawn from the prior distribution through a series of ABC-REJ iterations with decreasing tolerances $\epsilon_{1}>\dots>\epsilon_{T}$, where the posterior distribution from iteration $t-1$ becomes the prior distribution for iteration $t$. To encourage particle dispersion, the particles for iterations $t>1$ are perturbed with a Markov kernel $K_{t}(\theta|\theta^{\prime})$. The final set of particles $\left\{\theta_{T}^{(1)},\dots,\theta_{T}^{(N)}\right\}$ from iteration $T$ approximates the ABC-SMC posterior distribution $\pi_{ABC-SMC}\left(\theta\middle|d\left(s^{*},s^{obs}\right)<\epsilon_{T}\right)$. Note that $\pi_{ABC-SMC}\left(\theta\middle|d\left(s^{*},s^{obs}\right)<\epsilon_{1}\right)$ corresponds to $\pi_{ABC-rej}\left(\theta\middle|d\left(s^{*},s^{obs}\right)<\epsilon_{1}\right)$ from the ABC-REJ method.
1 for $t=1,\ldots,T$ do
2       for $i=1,\dots,N$ do
3             if $t=1$ then
4                  Sample $\theta^{**}\sim\pi(\theta)$
5            else
6                   Sample $\theta^{*}$ from $\left\{\theta^{(i)}_{t-1}\right\}$ with weights $\left\{w^{(i)}_{t-1}\right\}$
7                  $\theta^{**}\sim K_{t}(\theta|\theta^{*})$
8                  If $\pi(\theta^{**})=0$, then return to Step 6
9
10            Simulate data $y^{**}$ with parameter $\theta^{**}$, and compute $s^{**}=S(y^{**})$
11            If $d(s^{**},s^{obs})\geq\epsilon_{t}$, then return to Step 3
12            Record $\theta_{t}^{(i)}\leftarrow\theta^{**}$ with weight $w_{t}^{(i)}=\begin{cases}1&\text{ if }t=0\\ \frac{\pi\left(\theta^{(i)}_{t}\right)}{\sum^{N}_{j-1}w_{t-1}^{(j)}\cdot K_{t}\left(\theta_{t-1}^{(j)}\middle|\theta_{t}^{(i)}\right)}&\text{ if }t>0\end{cases}$
13      Normalize the weights such that $\sum_{i=1}^{N}w_{t}^{(i)}=1$
Algorithm 10 ABC-SMC